OpenStudy (anonymous):

solve 5+square root x+2 = 8+ square root x-7 note; the x+2 and x-7 are under the square root sign

hartnn (hartnn):

subtract 5 from both sides first, what do u get ?

OpenStudy (anonymous):

\[\sqrt{x+2} = 3 + \sqrt{x-7}\]

hartnn (hartnn):

now square both sides! notice that one of the square root (on left) gets eliminated! :)

OpenStudy (anonymous):

so my answer is \[3 + \sqrt{x-7}\]?

OpenStudy (anonymous):

wait.. if that's squared it'd 3+x-7

hartnn (hartnn):

no....the entire left side does not get eliminated, only the square root part, what do u get after you square both sides of that ?

OpenStudy (anonymous):

x+2=3+x-7?

hartnn (hartnn):

on the right side, the form is a+b and its square will be a^2+2ab+b^2 so now whats the square of 3 + sqrt (x-7) ??

OpenStudy (anonymous):

3^2+x-7

hartnn (hartnn):

nopes, there will be 3 terms a^2 b^2 2ab here a = 3 b=sqrt(x-7)

OpenStudy (anonymous):

okay so.. it'll be \[3^{2}+\sqrt{x-7}^{2}+2(3)(\sqrt{x-7})

OpenStudy (anonymous):

that didn't work very well... \[3^{2}+\sqrt{x-7}^{2}+2(3)(\sqrt{x-7})\]

hartnn (hartnn):

but thats correct, you just need to simplify the terms.

OpenStudy (anonymous):

\[9+\sqrt{x ^{2}-7^{2}} + \sqrt{6x-42}\] Like that?

hartnn (hartnn):

note \(\sqrt {stuff^2}=stuff\)

hartnn (hartnn):

and last term is just \(6 \sqrt{x-7} \)

OpenStudy (anonymous):

\[9+x-7 +6\sqrt{x-7}\]

hartnn (hartnn):

yes, simplify that

OpenStudy (anonymous):

\[2+x+6\sqrt{x-7}\] can whats in the sqrt simplify?

hartnn (hartnn):

and on the left side we had x+2 so, try to isolate sqrt (x-7) from \(x+2=2+x+6\sqrt{x-7}\)

hartnn (hartnn):

still there ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

goo , so what gets cancelled?

hartnn (hartnn):

**good

OpenStudy (anonymous):

the 9 and -7 get combined to 2

hartnn (hartnn):

we are simplifying this, right ? \(x+2=2+x+6\sqrt{x-7}\) the x+2 gets cancelled, what remains ?

OpenStudy (anonymous):

only the \[6\sqrt{x-7}\] is left isn't it?

hartnn (hartnn):

yes! but on the left side there's a 0 so, 0=6 sqrt {x-7} square both sides, what do u get ?

OpenStudy (anonymous):

you can't change 0 by squaring it so \[0 = 6(x-7)\] is that what you get?

hartnn (hartnn):

0 = 36 (x-7) but the point is, that x-7 = 0 !! got this ? so what's x ?

hartnn (hartnn):

it would be easy to figure out that x-7=0 gives you x=7 ask if any doubts in any steps :)

OpenStudy (anonymous):

oh okay so my final answer is x=7?

hartnn (hartnn):

yes! :)

hartnn (hartnn):

you can verify it by plugging in x=7 in your original equation...

OpenStudy (anonymous):

Awesome! Thank you very much for helping me! I'm not good at math at all.. so I always appreciate it when people take the time to help me :)

hartnn (hartnn):

you are most welcome ^_^ with enough practice, you can become better at math than me !! :P good luck! :)

OpenStudy (anonymous):

Thanks again :)